OpenStudy (anonymous):

find the Riemann sum for f(x)=x^2+x over the interval(0,8),where x0=0,x1=1,X2=3,X3=7,X4=8,and where c1=1,c2=2,c3=5 and c4=8

OpenStudy (loser66):

What are c's??

OpenStudy (loser66):

center??

OpenStudy (loser66):

\(\sum_{[x_n-x_{n-1}]} f(x) dx\) now, at c1= 1, f(1) = 1^2 +1 =2 and \(\triangle x= 1-0 =1\) hence, the first term is 2*1 the second term: at c2= 2, f(2) =2^2+2 = 6 and \(\triangle x = 3-1=2\) hence the second term is 6*2 =12 do the same with the third term and the fourth term add them together. done

OpenStudy (loser66):